OpenStudy (anonymous):

Determine the normalization constants for the following functions, given the form of the volume element dV and the limits of the variables psi(x)=x exp(-x^2/2) , dV=dx , -infinity

OpenStudy (anonymous):

The question means that you have to choose a normalisation constant, let's call it k, such that \[k \int\limits_{- \infty}^{+\infty} x e^{-\frac{ x^2 }{ 2 }}dx = 1\] So to find k you just need to evaluate the integral

OpenStudy (anonymous):

The question doesn't make sense to me, looking again, because the function given is an odd function, ( f(-x) = -f(x) ), so the integral from -infinity to +infinity will be zero. Are you sure about the limits ? 0 to infinity would have made sense.